in the harood domar model
in the harrod domar model ,if a countrygdp is constant at the smae time its saving rate is 30% and depreciation rate is 10% what muat be incremental cpaital outpur ratio
according to Maddison betwwen year 1 amd 1000BC world gdp percapita grew at an annual average rate of
Answer #1

Growth rate = saving rate/ capital output ratio - depriciation rate

=> g = s/ c - d

=> g = 0.3/c - 0.1

=> c(g+ 0.1) = 0.3

=> c = 0.3/(g+0.1)

Where c is incremental capital output ratio, we can put the value of growth rate(g) here to get the value of c.